40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 2259     Entry time: Thu Nov 12 17:24:29 2009
Author: Koji, Joe, Peter 
Type: Configuration 
Category: CDS 
Subject: ETMY CDS test started 

We started the test of the new CDS system at ETMY.

The plan is as follows:
We do the ETMY test from 9:30 to 15:00 at ETMY from Nov 12~17. This disables the ETMY during this period.
From 15:00 of the each day, we restore the ETMY configuration and confirm the ETMY work properly.


Today we connected megatron to the existing AA/AI modules via designated I/F boxes. The status of the test was already reported by the other entry.

During the test, c1iscey was kept running. We disabled the ETMY actuation by WatchDog. We did not touch the RFM network.

After the test we disconnected our cables and restored the connection to ICS110B and the AI/AA boards.

The WatchDog switches were released.

The lock of the ETMY was confirmed. The full interferometer was aligned one by one. Left in the full configuration with LA=off.

ELOG V3.1.3-